MM with Corporate Taxes

Companies U and L are identical in every respect except that U is unlevered while L has $16 million of 8% bonds outstanding. Assume: (1) All of the MM assumptions are met. (2) Both firms are subject to a 25% federal-plus-state corporate tax rate. (3) EBIT is $4 million. (4) The unlevered cost of equity is 12%.

  1. What value would MM now estimate for each firm? (Hint: Use Proposition I.) Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55. Round your answers to two decimal places.

    Company U: $   million
    Company L: $   million

  2. What is rs for Firm U? Round your answer to one decimal place.

    %

    What is rs for Firm L? Do not round intermediate calculations. Round your answer to one decimal place.

    %

  3. Find SL, and then show that SL + D = VL results in the same value as obtained in Part a. Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55. Do not round intermediate calculations. Round your answers to two decimal places.

    SL = $   million

    SL + D = $   million

  4. What is the WACC for Firm U? Do not round intermediate calculations. Round your answer to two decimal places.

    %

    What is the WACC for Firm L? Do not round intermediate calculations. Round your answer to two decimal places.

    %
Answer #1

Solution- Given data, Companies and ho are identical in every respect are except that is unlevered while thas $16 million of. (b) fs for firm o, &su - Unlevered cost of equity -129. rs for firm L, VSL = rsut Cosu-86) *(1-tax rateje bear steck = 127.ralue of firm L = 14.16+16 = $30.16 million,